OpenStudy (mhchen):

Non-Linear Mapping Transformations: f(x) = |x| If we apply the mapping: T(x,y) = (e^x , y) What will be the new equation?

OpenStudy (mhchen):

\[f(x)=|x|\\T(x,y) = (e^x,y)\]

OpenStudy (caozeyuan):

its kind of confusing since the first funciton maps R to R but the seocnd maps R^2 to R^2

OpenStudy (mhchen):

I think it's like \[f(x) = |ln{x}|\] Since that's the inverse....

OpenStudy (caozeyuan):

TBH, Im not sure what your trying to do here, can you explain the question alittle bit more

OpenStudy (mhchen):

So we have f(x) = |x| as our parent function. I wanted to know what the new function would be after the mapping of \[T(x,y)=(e^{x},y)\]

OpenStudy (caozeyuan):

so you what to know what si T(f(x))?

OpenStudy (caozeyuan):

wouldnt it be (e^abs(x),y)?

OpenStudy (mhchen):

T(x,y) is mapping notation. I'm pretty sure that for the x, you would need to find the inverse of it.

OpenStudy (caozeyuan):

maybe my brain is not working right now, but I am not understnaindg anything

OpenStudy (caozeyuan):

so what is this thing that you need? is it T(f(x)) or f(T(x)) or something else?

OpenStudy (caozeyuan):

f(T(x)) doesnt make sense becuase output of T is in R^2 and input of f is in R

OpenStudy (mhchen):

like if we have a function f(x) = x And we apply T(x,y) = (3x,y+1) The new function would be f(x) = (x/3) + 1

OpenStudy (mhchen):

That's what a mapping notation is essentially.

OpenStudy (mhchen):

T(x,y) is not a function, it's a mapping notation

OpenStudy (caozeyuan):

I totally dont see a difference between mapping and function

OpenStudy (caozeyuan):

sorry Im a biochemist not mathmatician so a lot of math I do everything are just routine calculation

OpenStudy (steve816):

yayy you finally got a profile pic mhchen!

OpenStudy (mhchen):

It's Elsa and leenathan's cult

OpenStudy (mhchen):

okay whatever, thanks for trying though. I think I got it myself, I can't stay on this question too long because I have a ton of others to review

OpenStudy (caozeyuan):

I still dont quite understand this mapping thngy, so Im gonna fromulate in my own language and you tell me if it is right

OpenStudy (caozeyuan):

T(x,y)=(f(x),g(y)), then the function you want is g( f^-1(x))

OpenStudy (caozeyuan):

thats why you take 3x and turned it to x/3 then add 1